OpenStudy (anonymous):

What is the solution to p over zero point three – 2 = 6?

OpenStudy (jadeishere):

\[\frac{ p }{ 0.3 } - 2 = 6\] Is that the equation?

OpenStudy (anonymous):

Yup!

OpenStudy (jadeishere):

kay, do you understand the rule that whatever you do to one side, you must do the to the other?

OpenStudy (anonymous):

I don't understand it at all. If you could explain it that would be great. (:

OpenStudy (jadeishere):

Kaykay :) so the rule I told you (what you do to one side of the equality sign, you MUST do to the other) works for every equation when you're goal is to isolate a variable. In this case, we're trying to get every number over to the RIGHT side of the equality (=) sign, to isolate 'p'. 1) \[\frac{ p }{ 0.3 } - 2 + (2) = 6 + (2)\] See what I did? I added 2 to -2 to eliminate 2 from the left side of the (=) sign, then added it to the right side as well.

OpenStudy (anonymous):

That easy? :o Oh my gosh. Lol. Thank you so much! I can't believe I couldn't understand that.

OpenStudy (jadeishere):

Lol yeah, it's easy once you understand it :) Steps we've done; Step 1; Simplify both sides Step 2; add 2 to both sides. The next step is to divide both sides by 3.3333335

OpenStudy (jadeishere):

so then 3.33333333p / 3.3333333335 = 8/3333333333.5 so, if you usa a calculator to get the answer (I'm not the best at doing this part by hand, you'll get p = 2.4 I hope this helps :)

OpenStudy (anonymous):

Okay! Thank you so much. People just tell me the answer, but you actually helped. Thanks.

OpenStudy (jadeishere):

Yep :)